# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

### v1

@@ Abstract
Low-rank adapters have become a standard approach for efficiently fine-tuning large language models (LLMs), but they often fall short of achieving the performance of full fine-tuning. We propose a method, LoRA Silver Bullet or LoRA-SB, that approximates full fine-tuning within low-rank subspaces using a carefully designed initialization strategy. We theoretically demonstrate that the architecture of LoRA-XS—which inserts a trainable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—provides the precise conditions needed for this approximation. We leverage its constrained update space to achieve optimal scaling for high-rank gradient updates while removing the need for hyperparameter tuning. We prove that our initialization offers an optimal low-rank approximation of the initial gradient and preserves update directions throughout training. Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of standard LoRA while using 27-90x fewer parameters, and comprehensively outperforms LoRA-XS. Our findings establish that it is possible to simulate full fine-tuning in low-rank subspaces, and achieve significant efficiency gains without sacrificing performance. We have released our code publicly at https://github.com/RaghavSinghal10/lora-sb.
@@ 1 Introduction
Pre-trained language models have become central to natural language processing, achieving state-of-the-art performance across diverse tasks (Radford et al., 2021; Kirillov et al., 2023; Achiam et al., 2023). While these models excel at general-purpose capabilities (Bubeck et al., 2023; Hao et al., 2022), adapting them to specific downstream tasks often requires fine-tuning. Although in-context learning (Brown et al., 2020; Radford et al., 2019) has gained popularity for its simplicity, it falls short in both performance and efficiency compared to fine-tuning (Liu et al., 2022). At the same time, full fine-tuning, while highly effective, is computationally expensive and impractical at scale, highlighting the need for more efficient adaptation techniques.
Parameter-efficient fine-tuning (PEFT) has emerged as a crucial approach for adapting large language models (LLMs) while addressing computational constraints. While full fine-tuning (FFT) typically achieves optimal performance, it requires updating billions of parameters, making it computationally intensive and often impractical for most applications. Low-rank decomposition methods, beginning with LoRA (Hu et al., 2021), have shown particular promise by significantly reducing trainable parameters through learning low-rank updates. This has sparked numerous advances in low-rank methods that either enhance performance through better optimization techniques and initialization strategies, or improve parameter efficiency through structured matrices and adaptive rank selection (Zhang et al., 2023; Wang et al., 2024b, a). However, these methods still face fundamental trade-offs, they must either maintain a relatively large number of parameters to match FFT performance, or accept a performance degradation when pursuing extreme parameter efficiency (Hu et al., 2021; Ding et al., 2023; Wang et al., 2024b). This raises an important question: can we design low-rank methods that maintain FFT-competitive performance while drastically reducing the parameter count beyond current approaches?
@@
Figure 1: LoRA-SB Illustration. LoRA-XS (Bałazy et al., 2024) reduces parameter count compared to LoRA (Hu et al., 2021) by inserting a trainable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$. Our method, LoRA-SB, leverages the same architecture. We find that updating $R$ using its gradients $g^{R}$ is equivalent to updating the full-finetuning matrix $W$ with an equivalent gradient $\tilde{g}_{SB}=sBg^{R}A$. We initialize $B$, $R$, and AA such that the equivalent gradient $\tilde{g}_{SB}$ optimally approximates the full fine-tuning gradient $g$ in low rank subspaces at each training step. In essence, we simulate the entire full fine-tuning process optimally within low-rank subspaces by utilizing only the first gradient $g_{1}$ (shown in green) from full fine-tuning.
@@
Low-rank decomposition methods operate on a fundamental premise: fine-tuning requires learning only a low-rank update to the pre-trained weights. Some theoretical work extends this hypothesis, suggesting that methods like LoRA can learn any low-rank approximation of the full fine-tuning gradient. However, the gradients computed by these methods to update their trainable adapters do not inherently possess this property. For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full fine-tuning gradient (Wang et al., 2024b). Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like Pissa-LoRA (Meng et al., 2024) and LoRA-GA (Wang et al., 2024a), which address this challenge through various low-rank approximation techniques.
@@
We formally analyze these limitations in the context of the architecture used in LoRA-XS (Bałazy et al., 2024)—which inserts a trainable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—and demonstrate that these challenges are even more pronounced in its framework. While exploring solutions inspired by existing LoRA-based methods, we discover a remarkable property unique to LoRA-XS: through careful initialization of matrices $A$ and $B$, we can approximately simulate the full fine-tuning optimization in low rank subspaces throughout the entire training, as shown in Figure 1. Our initialization strategy provides optimal scaling for approximating high-rank full fine-tuning gradients and eliminates the need for the scaling hyperparameter $\alpha$, results which we prove theoretically. Our key contributions include:
@@
  • We propose a principled initialization strategy derived from approximating the first step of full fine-tuning, proving it provides optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
@@
  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method surpasses the performance of LoRA while using 27-90x less parameters, and comprehensively outperforms LoRA-XS.
@@ 2 Related Work
Parameter-Efficient Fine-Tuning. Parameter-Efficient Fine-Tuning (PEFT) methods have become essential for adapting large pre-trained models under computational constraints. Early techniques like AdapterFusion (Pfeiffer et al., 2021) and Prefix-Tuning (Li & Liang, 2021) enabled task-specific adaptation with minimal parameter updates. Advances like soft prompts (Lester et al., 2021) further reduced trainable parameter counts while maintaining strong performance. Recent approaches have explored operating directly on model representations (Wu et al., 2024), offering different trade-offs between efficiency and performance.
Low-Rank Decomposition Methods. LoRA (Hu et al., 2021) demonstrated that weight updates during fine-tuning could be efficiently approximated using low-rank matrices, drastically reducing parameter counts while freezing pre-trained weights. Building on this insight, variants such as QLoRA (Dettmers et al., 2023) and AdaLoRA (Zhang et al., 2023) extended the paradigm through quantization and adaptive allocation strategies. The applicability of low-rank techniques has also been explored in pretraining with GaLore (Zhao et al., 2024) and ReLoRA (Lialin et al., 2023), highlighting the versatility of low-rank adaptation in both fine-tuning and pre-training contexts. LoRA-based methods have also been applied in other domains, such as efficient federated fine-tuning (Sun et al., 2024; Singhal et al., 2024).
Enhancing LoRA Performance. Recent efforts have focused on optimizing LoRA’s performance. PiSSA-LoRA (Meng et al., 2024) demonstrated faster convergence and improved task performance by initializing matrices with principal components of pre-trained weights. LoRA-Pro (Wang et al., 2024b) and LoRA-GA (Wang et al., 2024a) improved gradient approximation, aligning low-rank updates more closely with full fine-tuning. Other methods like DoRA (Liu et al., 2024) and rsLoRA (Kalajdzievski, 2023) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
Improving Efficiency in LoRA Variants. Efficiency-focused innovations have pushed LoRA toward even greater parameter savings. LoRA-XS (Bałazy et al., 2024) achieves parameter reduction by inserting a small trainable weight matrix into frozen low-rank matrices. VeRA (Kopiczko et al., 2024) shares low-rank matrices across layers, relying on scaling vectors for task-specific adaptation. Tied-LoRA (Renduchintala et al., 2024) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (Tian et al., 2024) introduces an asymmetric architecture to improve adaptation in complex domains without domain-specific expertise.
@@ 3.1 Preliminaries
In standard fine-tuning, a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$ is updated as:
@@
where $W_{0}$ is the pre-trained weight. This requires updating $mn$ parameters per layer.
LoRA hypothesizes that these updates lie in a low-dimensional subspace and parameterizes the update matrix $\Delta W$ as:
@@
where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are trainable low-rank matrices, with rank $r\ll\min(m,n)$, and $s$ is a scaling factor to stabilize training. This reduces parameters from $mn$ to $r(m+n)$.
LoRA-XS introduces a more efficient parameterization:
@@
where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are fixed matrices, and only $R\in\mathbb{R}^{r\times r}$ is trainable, reducing parameters to $r^{2}$.
We denote full fine-tuning gradient as $g=\frac{\partial L}{\partial W}$, and LoRA-XS gradient as $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$.
@@ 3.2 Motivation
LoRA-XS (Bałazy et al., 2024), although having significantly fewer learnable parameters compared to LoRA, exhibits suboptimal performance. The difference in architecture causes constraints on the type of updates LoRA-XS can learn. The subspace of learned updates is characterized in Lemma 1.
@@
Let $\Delta W$ be an update learned by LoRA-XS . Then, the set of all $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, can be given by $\{M\in\mathbb{R}^{m\times n}|\text{Col}(W)\subseteq\text{Col}(M)$  $\wedge\text{ Row}(M)\subseteq\text{Row}(A)\}$
@@
Appendix A.2
The above lemma implies that while $\Delta W$ is constrained to be $r$ ranked or lower, it also needs to have row and column spaces defined by those of $B$ and $A$ respectively. As against this, LoRA can learn any update $\Delta W$ as long as $\text{rank}(\Delta W)\leq r$. Thus the low expressivity of LoRA-XS as compared to LoRA can account for the performance drop. We identify three key limitations which arise due to this and otherwise:
@@
1) Inadequate Gradient Approximation: LoRA optimization is mathematically equivalent to full fine-tuning using a constrained low-rank gradient. The gradient of LoRA does not optimally approximate the full gradient, and needs to be tuned at each step. LoRA-Pro (Wang et al., 2024b) finds that this results in suboptimal performances, and provides a closed form solution to optimize the gradients. In LoRA-XS , the gradient updates are restricted to an even more constrained low-rank space since the matrices $A$ and $B$ are fixed. We posit that the limitation becomes particularly severe when the ideal updates lie outside the space spanned by fixed $A$ and $B$, and consequently has a larger impact on performance.
@@
2) Suboptimal Initialization: While initialization impacts all low-rank methods, it becomes critical in LoRA-XS where $A$ and $B$ are frozen. Unlike LoRA where poor initialization can be compensated through training, LoRA-XS must rely entirely on its initial subspace defined by $A$ and $B$. Consider the zero initialization of the $B$ matrix, for example. While LoRA may experience some performance degradation in this case (Wang et al., 2024a) (Meng et al., 2024), the ideal low-rank update $\Delta W$ can still be reached through gradient descent. In fact, zero initialization for the $B$ matrix is a commonly used choice, including in the original LoRA implementation (Hu et al., 2021). However, in the case of LoRA-XS , this would result in no learning, as the product $BRA$ would remain zero. The current LoRA-XS method leverages the most significant subspaces spanned by the columns of pre-trained weights $W_{0}$, inspired by (Meng et al., 2024). This initialization is inadequate because it fails to capture the specific subspaces relevant to the fine-tuning task. This issue, known to affect LoRA ’s performance (Wang et al., 2024a), could have an even greater impact on LoRA-XS for the reasons discussed above.
@@
3) Hyperparameter Sensitivity: The scaling factor $\alpha$, present in almost every LoRA based fine-tuning method requires careful tuning to maintain stability during training. This hyperparameter acts as a bridge between the low-rank and full-rank spaces, compensating for the dimensional mismatch in gradients. Poor tuning of $\alpha$ can lead to unstable training or slow convergence (Kalajdzievski, 2023), adding complexity to the adaptation process and potentially limiting practical deployment.
3.3 Approximation of the full fine-tuning gradient
@@
As mentioned above, LoRA optimization is mathematically equivalent to full fine-tuning using a constrained low-rank gradient. However the update generated using the gradients of LoRA does not result in the same update which the low-rank gradient would have generated. The following holds true for LoRA-XS as well.
To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$. The relationship can be simply given by $dW=\frac{\partial W}{\partial R}g^{R}$. This relationship implies that updating matrix $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full fine-tuning as described in Definition 1.
Definition 1 (Equivalent Gradient).
In the context of LoRA-XS optimization, we define the equivalent gradient as: $\tilde{g}=sBg^{R}A$ where $s$ is the scaling factor, and $g^{R}$ is the gradient with respect to matrix $R$.
@@
The equivalent gradient describes the virtual low-rank gradient of matrix $W$ in LoRA-XS optimization process, despite $W$ not being directly trainable. This gradient determines how updates to $R$ affect the overall weight matrix. To bridge the performance gap between LoRA-XS and full fine-tuning, our goal is to minimize the discrepancy between the equivalent gradient $\tilde{g}$ and the full gradient $g$. First, we establish the relationship between gradients in LoRA-XS optimization in Lemma 2.
@@
During LoRA-XS optimization, the gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{T}gA^{T}$
@@
Appendix A.2
Leveraging this relationship, we can formulate our objective to minimize the distance between the equivalent gradient and the full gradient. We do not have access to the full fine-tuning gradient $g$ during LoRA-XS based fine-tuning. Thus we need to find the ideal gradient with respect to $R$, given by $g^{R}$, and subsequently the optimal approximation $\tilde{g}$, in terms of the gradient which is available to us during training: $g^{R}_{LoRA-XS}$. Fortunately, this optimization problem admits a closed-form solution independent of $g$ as described in Theorem 3.
@@
[…]
@@
Appendix A.3. ∎
[…]
@@
Consider the update for matrix $R$ using the closed-form solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta\geq 0$ is the learning rate. This update guarantees a reduction in the loss function, similar to traditional gradient descent methods. Specifically, the change in loss $\mathrm{d}L$ is given by: $\mathrm{d}L=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle\leq 0$

### v3

@@ Abstract
Low-rank adapters have become standard for efficiently fine-tuning large language models (LLMs), but they often fall short of achieving the performance of full fine-tuning. We propose a method, LoRA Silver Bullet or LoRA-SB, that approximates full fine-tuning within low-rank subspaces using a carefully designed initialization strategy. We theoretically demonstrate that the architecture of LoRA-XS—which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—provides the precise conditions needed for this approximation. We leverage its constrained update space to achieve optimal scaling for high-rank gradient updates while removing the need for hyperparameter tuning. We prove that our initialization offers an optimal low-rank approximation of the initial gradient and preserves update directions throughout training. Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of standard LoRA while using 27-90 times fewer learnable parameters, and comprehensively outperforms LoRA-XS. Our findings establish that it is possible to simulate full fine-tuning in low-rank subspaces, and achieve significant efficiency gains without sacrificing performance. Our code is publicly available at https://github.com/RaghavSinghal10/lora-sb.
@@ 1 Introduction
Pre-trained language models have become central to natural language processing, achieving state-of-the-art performance across diverse tasks (Radford et al., 2021; Kirillov et al., 2023; Achiam et al., 2023). While these models excel at general-purpose capabilities (Bubeck et al., 2023; Hao et al., 2022), adapting them to specific downstream tasks often requires fine-tuning (FT). Although in-context learning (Brown et al., 2020; Radford et al., 2019) has gained popularity for its simplicity, it falls short in both performance and efficiency compared to FT (Liu et al., 2022). At the same time, full FT, while highly effective, is computationally expensive and impractical at scale.
Parameter-efficient fine-tuning (PEFT) has become vital for adapting large language models (LLMs) under computational constraints. Low-rank methods like LoRA (Hu et al., 2021) address this by reducing learnable parameters via low-rank updates, sparking advancements in optimization, initialization, structured matrices, and adaptive rank selection (Zhang et al., 2023; Wang et al., 2024b, a). However, these methods face trade-offs: either retain many parameters to match full FT or sacrifice performance for extreme efficiency (Hu et al., 2021; Ding et al., 2023; Wang et al., 2024b). This raises a critical question: Can we design low-rank methods that achieve full FT-level performance while drastically reducing parameter counts?
@@
Figure 1: LoRA-SB. LoRA-XS (Bałazy et al., 2024) reduces parameters compared to LoRA (Hu et al., 2021) by inserting a learnable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$. Our method, LoRA-SB, uses the same architecture. We find that updating $R$ using its gradients $g^{R}$ is equivalent to updating the full FT matrix $W$ with an equivalent gradient $\tilde{g}_{SB}=sBg^{R}A$. We initialize $B$, $R$, and $A$ such that the equivalent gradient $\tilde{g}_{SB}$ provably best approximates the full FT gradient $g$ in low rank subspaces at each step. In essence, we simulate the entire full FT process optimally within low-rank subspaces by utilizing only the first gradient $g_{1}$ (shown in green) from full FT.
@@
Low-rank decomposition methods operate on a fundamental premise: FT requires learning only a low-rank update to the pre-trained weights. Some theoretical work extends this hypothesis, suggesting that methods like LoRA can learn any low-rank approximation of the full FT gradient. However, the gradients computed by these methods do not inherently possess this property. For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full FT gradient (Wang et al., 2024b). Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like Pissa-LoRA (Meng et al., 2024) and LoRA-GA (Wang et al., 2024a).
@@
We analyze these limitations in the context of the architecture of LoRA-XS (Bałazy et al., 2024)—which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—and demonstrate that these challenges are even more pronounced. While exploring solutions inspired by LoRA-based methods, we discover a remarkable property unique to LoRA-XS: through careful initialization of $A$ and $B$, we can simulate the full FT optimization in low rank subspaces through entire training, as shown in Figure 1. Our initialization provides optimal scaling for approximating high-rank full FT gradients and eliminates need for scaling the hyperparameter $\alpha$.
Our key contributions are:
@@
  • We propose an initialization strategy derived from approximating the first step of full FT, proving it provides optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
@@
  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method surpasses LoRA’s performance while using 27-90x less learnable parameters, and comprehensively outperforms LoRA-XS.
@@ 2 Related Work
Parameter-Efficient Fine-Tuning (PEFT). PEFT methods have become essential for adapting large pre-trained models under computational constraints. Early techniques like AdapterFusion (Pfeiffer et al., 2021) and Prefix-Tuning (Li & Liang, 2021) enabled task-specific adaptation with minimal parameter updates. Advances like soft prompts (Lester et al., 2021) further reduced trainable parameter counts while maintaining strong performance. Recent approaches have explored operating directly on model representations (Wu et al., 2024).
Low-Rank Decomposition Methods. LoRA (Hu et al., 2021) demonstrated that weight updates during FT could be efficiently approximated using low-rank matrices, drastically reducing parameter counts. Building on this insight, variants such as QLoRA (Dettmers et al., 2023) and AdaLoRA (Zhang et al., 2023) extended the paradigm through quantization and adaptive allocation strategies. The applicability of low-rank techniques has also been explored in pretraining with GaLore (Zhao et al., 2024) and ReLoRA (Lialin et al., 2023), highlighting the versatility of low-rank adaptation methods. LoRA-based methods have also been applied in other domains, such as efficient federated FT (Sun et al., 2024; Singhal et al., 2025).
Enhancing LoRA Performance. Recent efforts have focused on optimizing LoRA’s performance. PiSSA (Meng et al., 2024) demonstrated improvements by initializing matrices with principal components of pre-trained weights. LoRA-Pro (Wang et al., 2024b) and LoRA-GA (Wang et al., 2024a) improved gradient approximation, aligning low-rank updates more closely with full FT. Methods like DoRA (Liu et al., 2024) and rsLoRA (Kalajdzievski, 2023) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
Improving Efficiency in LoRA Variants. Efficiency-focused innovations have pushed LoRA toward more parameter savings. LoRA-XS (Bałazy et al., 2024) achieves this by inserting a small trainable weight matrix into frozen low-rank matrices. VeRA (Kopiczko et al., 2024) shares low-rank matrices across layers, relying on scaling vectors for task-specific adaptation. Tied-LoRA (Renduchintala et al., 2024) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (Tian et al., 2024) introduces an asymmetric architecture for improvement.
@@ 3.1 Preliminaries
In standard FT, a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$ is updated using the update matrix $\Delta W$ as:
@@
where $W_{0}$ is the pre-trained weight. This requires updating $mn$ parameters per layer. LoRA posits that updates lie in a low-dimensional subspace, parameterizing $\Delta W$ as:
@@
where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are trainable low-rank matrices with rank $r\ll\min(m,n)$, and $s$ is a scaling factor ($\alpha/r$) to stabilize training. This reduces the number of parameters from $mn$ to $r(m+n)$. LoRA-XS efficiently parameterizes as:
@@
where $B$ and $A$ are fixed, and only $R\in\mathbb{R}^{r\times r}$ is trainable, reducing the number of parameters to $r^{2}$. We denote the full FT gradient as $g=\frac{\partial L}{\partial W}$, and the LoRA-XS gradient as $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$, where $L$ is the loss function.
@@ 3.2 Motivation
LoRA-XS (Bałazy et al., 2024) has significantly fewer learnable parameters than LoRA but performs suboptimally. LoRA-XS’s architecture causes constraints on the type of updates it can learn. The subspace of learned updates is characterized in Lemma 1. This implies that while $\Delta W$ is constrained to be rank $\leq r$, it also needs to have column and row spaces defined by those of $B$ and $A$, respectively. In contrast, LoRA can learn any update $\Delta W$ as long as $\text{rank}(\Delta W)\leq r$. Thus, the low expressivity of LoRA-XS as compared to LoRA can account for the performance drop.
@@
Let $\Delta W$ be an update learned with LoRA-XS. Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)$ $\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where $\text{Col}(M)$ and $\text{Row}(M)$ are column and row spaces of matrix $M$ respectively.
@@
See Appendix A.1. ∎
We identify three key limitations, which arise due to this and otherwise:
@@
1) Inadequate Gradient Approximation: LoRA optimization is mathematically equivalent to full FT using a constrained low-rank gradient. The gradient of LoRA does not optimally approximate the full gradient, and needs to be tuned at each step. LoRA-Pro (Wang et al., 2024b) finds that this results in suboptimal performances, and provides a closed form solution to optimize the gradients. In LoRA-XS, the gradient updates are restricted to an even more constrained low-rank space since $A$ and $B$ are fixed. We posit that the limitation becomes particularly severe when the ideal updates lie outside the space spanned by fixed $A$ and $B$, and consequently has a larger impact on performance.
@@
2) Suboptimal Initialization: While initialization impacts all low-rank methods, it becomes critical in LoRA-XS where $A$ and $B$ are frozen. Unlike LoRA where poor initialization can be compensated through training, LoRA-XS relies entirely on its initial subspace defined by $A$ and $B$. Consider the zero initialization of the $B$ matrix, for example. While LoRA may experience some performance degradation in this case (Wang et al., 2024a; Meng et al., 2024), the ideal low-rank update $\Delta W$ can still be reached through gradient descent. In fact, zero initialization for the $B$ matrix is commonly used, including in the original LoRA paper (Hu et al., 2021). However, in LoRA-XS, this results in no learning, as the product $BRA$ remains zero. LoRA-XS uses the most significant subspaces spanned by the columns of pre-trained weights for initialization, inspired by Meng et al. (2024). This initialization is not aligned well with FT because it fails to capture the specific subspaces relevant to the FT task.
@@
3) Hyperparameter Sensitivity: The scaling factor $s$, present in almost every LoRA based FT method requires tuning to maintain stability during training. This factor acts as a bridge between the low-rank and full-rank spaces, compensating for the dimensional mismatch in gradients. Poor tuning of $s$ can lead to unstable training or slow convergence, e.g., see rsLoRA (Kalajdzievski, 2023), adding complexity and potentially limiting practical deployment.
3.3 Approximation of the full FT gradient
@@
As mentioned, LoRA optimization is mathematically equivalent to full FT using a constrained low-rank gradient. However, the update generated using the gradients of LoRA does not result in the same update which the low-rank gradient would have generated. The following holds true for LoRA-XS as well. To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$, which can be simply given by $\mathrm{d}W=-sB(\mathrm{d}R)A.$ This implies that updating $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full FT as described in Definition 1.
Definition 1.
We define the equivalent gradient as: $\tilde{g}=sBg^{R}A$ where $g^{R}$ is the gradient of $L$ with respect to $R$.
@@
The equivalent gradient describes the virtual low-rank gradient of matrix $W$ in LoRA-XS optimization process, despite $W$ not being directly trainable. This gradient determines how updates to $R$ affect $W$. To bridge the performance gap between LoRA-XS and full FT, we aim to minimize the discrepancy between the equivalent gradient $\tilde{g}$ and the full gradient $g$. First, we establish the relationship between gradients in LoRA-XS optimization in Lemma 2.
@@
The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$
@@
See Appendix A.2. ∎
We can now formulate our objective to minimize the distance between the equivalent gradient and the full gradient. We do not have access to the full FT gradient $g$ during LoRA-XS based FT. Thus we need to find the ideal gradient with respect to $R$, given by $g^{R}$, and subsequently the optimal approximation $\tilde{g}$, in terms of the gradient which is available to us during training: $g^{R}_{LoRA-XS}$. Fortunately, this optimization problem admits a closed-form solution independent of $g$ as described in Theorem 3.
@@
The optimal solution for the objective $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, such that $\tilde{g}=sBg^{R}A$, is: $g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}$ (4)
@@
See Appendix A.3. ∎
The closed-form solution in Theorem 3 solves the optimization problem $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, but by itself doesn’t ensure the loss will decrease when updating $R$. Through Theorem 4, we prove that the change in loss is non-positive ($\Delta L\leq 0$). This property is fundamental to optimization as it guarantees consistent loss minimization throughout training.
@@
Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate. This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0.$